What is the equation of a line that passes through (4, 6) and (6, 12)?
Answer by mananth(16949)

x1 y1 x2 y2
4 6 6 12

slope m = (y2-y1)/(x2-x1)
( 12 - 6 )/( 6 - 4 )
( 6 / 2 )
m= 3.00 3

Plug value of the slope and point ( 4,6) in
Y = m x + b
6.00 = 12 + b
b= 6 -12
b= -6
So the equation will be
Y = 3 x -6
m.ananth@hotmail.ca